OpenStudy (anonymous):

I don't know how to start this one. (t/t-4)=(t+4/6)-1

OpenStudy (anonymous):

@Precal

OpenStudy (anonymous):

\[(\frac{ t }{ t-4 }) = (\frac{ t+4 }{ 6 }) -1\] can you solve proportions?

OpenStudy (anonymous):

do you think that is how you would go about solving it? or what did you learn in school about doing this?

OpenStudy (anonymous):

you actually can simplify these terms like t/t = 1

OpenStudy (anonymous):

No I can not solve proportions

OpenStudy (anonymous):

so for the first one you would have -1/4 = ... can you do the other one?

OpenStudy (anonymous):

-1/4 = (t+2/3) -1 lets multiply both sides by -1 to get rid of it 1/4 =-(t+2/3)

OpenStudy (anonymous):

Where did you get t/t

OpenStudy (anonymous):

t/t-4 t/t is 1

OpenStudy (anonymous):

\[(\frac{ t }{ t-4 }) = (\frac{ t+4 }{ 6 }) -6 \] \[(\frac{ t }{ t-4 }) = \left( \frac{ t+4-1 }{ 6 } \right)\] \[(\frac{ t }{ t-4 }) - \left( \frac{ t+4-1 }{ 6 } \right) \]

OpenStudy (anonymous):

-2/3 both sides you get 3/12 - 8/12 -5/12 = -t t= 5/12

OpenStudy (anonymous):

\[\frac{ \left( 6 \times t \right) - \left( t-4 \right)\left( t+3 \right) }{ 6\left( t-4 \right) }\]

OpenStudy (anonymous):

I dont even know if that is right cuz mites is doing something diff. i would guess he is right cuz i've only done Algebra I. But i just tried to help you from what i know.

OpenStudy (anonymous):

r.h.s =0 \[6t - \left( t^2 - t - 12 \right) = 0\]

OpenStudy (anonymous):

I am sorry the actual problem has no parenthacies. I don't not know how to write it

OpenStudy (anonymous):

\[6t - t^2 +t +12 = 0\] \[t^2 - 7t - 12 = 0\] \[\left( t-4 \right)\]

OpenStudy (anonymous):

sry \[\left( t-4 \right)\left( t-3 \right)\]

OpenStudy (anonymous):

therefore t=4 or t =3

OpenStudy (anonymous):

\[(\frac{ t }{ t-4 }) = (\frac{ t+4 }{ 6 }) - \frac{ 6 }{ 6 }=0\] \[(\frac{ t }{ t-4 }) = \left( \frac{ t+4-6 }{ 6 } \right) = 0 \] \[(\frac{ t }{ t-4 }) - \left( \frac{ t+4-6 }{ 6 } \right) = 0\] \[\frac{ \left( 6 \times t \right) - \left( t-4 \right)\left( t-2 \right) }{ 6\left( t-4 \right) }=0\] \[6t - \left( t^2 - 6t +8 \right) = 0\] now you can solve